OpenStudy (anonymous):

in a triangle,prove that sin(c/2)=cos((a+b)/2) (doesnt matter what triangle)

OpenStudy (raden):

a+b+c=180, so c = 180 - (a+b) this is a start

OpenStudy (anonymous):

allrdy there :p

OpenStudy (raden):

c = 180 - (a+b) divided by 2 on both sides, get c/2 = 90 - (a+b)/2

OpenStudy (raden):

plug sin on both sides, get sin(c/2) = sin(90 - (a+b)/2)

OpenStudy (raden):

finally, use the identity : sin(90 - x) = cosx, to simplify the right side above. get sin(c/2) = sin(90 - (a+b)/2) sin(c/2) = cos((a+b)/2)

OpenStudy (raden):

does that make sense ?

OpenStudy (anonymous):

ye really thx dude :p

OpenStudy (raden):

alhamdulillah, you're welcome bro :)